# NeRF-Aug: Data Augmentation for Robotics with Neural Radiance Fields

### v2

@@ Abstract
I Introduction
@@
Figure 1: When a human provides expert demonstrations for training a behavior cloning model, the model is effective for the object in the demonstration, but will fail for novel objects. We propose NeRF-Aug, where we automatically learn NeRFs for the novel objects and inpaint them in the expert data. With this photorealistic synthetic data, the robot can learn to interact successfully with novel objects.
@@
Figure 2: Comparison of inpainting results when replacing the original object with various objects in our training set. Instruct Pix-to-Pix, GenAug, and RoboAgent create images that show a new object, but do not recreate our specific object with appropriate dimensions and texture, this generally causes these methods to fail. Dreambooth has trouble with inpainting complex masks and also fails to recreate exact shade of colors.
@@
  • We edit videos of existing demonstrations by removing the training object via in-painting and blend the NeRF render of a novel object into the inpainted image to generate a synthetic dataset.
  • We demonstrate effective generalization on five diverse tasks using the generated synthetic data for training.
II Related Works
II-A Imitation Leaning
@@
II-B Neural Radiance Fields (NeRFs)
In recent years, there has been a surge in research on NeRFs. [20, 21, 22, 23, 24] are all NeRF approaches that are capable of rendering high resolution images in real time. Other methods focus on enhancing the robustness of NeRF models against real-world constraints, such as a limited number of images [25], camera pose noise [26], and glare [27]. Finally, other novel-view synthesis techniques [28, 29, 30, 31] remove the traditional multi-layer perceptron used to volume render images with other trainable structures.
Within the context of robotics, NeRFs have been used for robot navigation by creating a 3D map of the environment  [32][33, 34] combine NeRFs and reinforcement learning.  [35, 36] explored using robot arms for creating higher quality NeRF models of objects. Conversely, several works have explored using NeRF models for robot decision making, including [37, 38] which creates accurate depth maps for transparent images using NeRF models, and [39] which creates an object level NeRF, imagines a scene with the object in a different place and queries a vision-language model (VLM) for feedback. Finally, [40] adds noisy viewpoints along a demonstration trajectory and take corrective actions to make a more robust behavior cloning model, and  [41, 42] combine structure from motion and NeRFs.
II-C Synthetic Data in Robotics
@@
III Preliminaries
Imitation Learning. In imitation learning, we assume a premade list of $N$ expert trajectories $D=\left\{\tau_{i}\right\}_{i=1}^{N}$ where each trajectory consists of state-action pairs $\tau_{i}=\{(s_{k},a_{k})\}_{k=1}^{K}$. In visual imitation learning, we assume access to $n$ cameras where each contributes an image to the state. Thus, our state can be denoted as $s_{k}=(I_{1},I_{2},...,I_{n})$ where $I_{k}$ is an image. In vanilla behavior cloning, the policy $\pi$ is trained offline and learns a mapping between states and actions from our expert dataset by minimizing the loss $\mathcal{L}(\theta)=\mathbb{E}_{(s,a)\sim D}\left[\mathscr{l}\left(\pi(s),a\right)\right]$ for a given distance metric $\mathscr{l}$.
@@ IV Our Approach: NeRF-Aug
IV-A Creating a NeRF model of the Novel Object
Given a novel object for which there exists no training data, we train a NeRF model for this object. Using a gripper mounted camera, we are able to collect a dataset of image-pose pairs $\left\{(I_{k},T_{k})\right\}_{k=1}^{M}$ by moving the gripper to various viewpoints both close and far away from the novel object. We use these images to train a NeRF model for the object. In our experiments, we used NerfStudio’s [50] default Nerfacto model, which refines the camera intrinsic and extrinsic parameters of all images to denoise measurements.
IV-B Calculating Camera Pose Relative to Object
@@
Next, NeRF-Aug finds the camera position relative to the object. Using the gripper position with respect to the world given by $T_{\text{gripper}}(t)\in\text{SE}(3)$, we are able to find the camera-to-world matrix of the gripper camera. We multiply the position of the gripper by the offset of the camera from the gripper center
@@
If the object is already grasped, the relative position of the camera and the object stays the same, denoted by
@@
We also add a small amount of noise at each timestep into the camera-to-object matrix before rendering in the NeRF model. This slightly shifts the object from its’ original position to simulate a larger degree of grasp positions.
Figure 4: Our setup. We have a Realsense D455 cameras, The robot arm is a UR5e from Universal Robotics. During our tests, we use a large number of background objects such as the two dish racks and the wooden chest to reflect a real world scenario.
Figure 5: A top-down view of the objects used in our real-world study. We used a variety of common household items of varying shapes, sizes and colors to show several kinds of generalization. Left are the novel objects, right are the original objects.
IV-C NeRF Renderings
@@
IV-D Combining NeRF Renderings and Original Image
@@
This results in a frame where the original object is removed from the frame and only the background objects/scene remain. From there, we blend the object pixels from the NeRF renderings onto the image where the object was removed. We combine $I_{\text{NeRF}}$ and $I_{\text{no-object}}$ to get a new image with the novel object at the same pose as the original object while still keeping the background of the original image. This is done by computing
@@ V Experiments
In our real-world experiments, we use a Universal Robotics UR5e robot arm with a Robotiq-85 gripper. The arm is equipped with a Realsense D455 mounted to the front of the gripper (see Fig. 4). We run all experiments on a single Nvidia RTX A4000 GPU. We use the default Nerfacto model from Nerfstudio [50] for training without any hyperparameter tuning aside from stopping the training at 4,000 timesteps for object NeRFs and 20,000 timesteps for background NeRFs.
@@
TABLE I: Task success rates for novel objects ($\uparrow$)
@@ V-A Baselines
V-B Real World Tasks
For the Remove Object task, our initial object is a yellow spatula that rests on top of a small frying pan. The robot grasps the spatula and lifts it from the pan then drops it to the left of the pan. For this task we test on 3 novel objects: a pair of blue tongs, a red spoon, and a long wooden stick. These objects show how well NeRF-Aug can generalize to different colors and lengths of objects. For this task, we used 30 demonstrations.
For the Place in Chest task, our initial object is a mustard bottle which the robot grasps and places into a wooden chest. We tested on two novel objects: a swiss miss cardboard box and a paperback book. These novel objects differ in geometry and texture from the original object. For this task, we used 30 demonstrations.
@@ V-B Real World Tasks
TABLE II: Time To Create New Data in Minutes Including NeRF Model Training ($\downarrow$)
@@
V-C Comparison to Baselines
As seen in Table I, our method consistently outperforms other methods on all five tasks. Over the next best method, we achieve an average of 43% increase in the ”Remove Objects” task, a 5% increase in the ”Place In Chest” task, an 80% increase on the ”Cook with Pot” task, and a 50% increase on the ”Lift up High” task. Our method achieved a 85.6% success rate compared to the next best method (Dreambooth) with a 30% success rate.
Which methods synthesize object locations and sizes accurately? GenAug, Instruct Pix-to-Pix, and RoboAgent often output images of an object in either the wrong orientation or position. For example, as seen in fig.  2, we prompted the baselines to inpaint a ”red mug,” only for the output image to consist of a sideways view of a mug that is much larger than the actual mug. In contrast, our method was able to accurately synthesize the red mug from a top-down view. We argue that being able to render an object at unconventional viewpoints, such as when an object is very close to the camera or in a birds-eye-view, is a major reason why our method performs better than the baselines (such as being able to lift the mug 90% more often than any other method). Dreambooth seems more effective at synthesizing the correct object viewpoint, but will fail when the original object is significantly different in shape from the old object (see fig. 2 where Dreambooth fails to inpaint a swiss miss box and a book when the original object was a mustard bottle). A wide degree of viewpoints is important because images where the object is very close to the camera or from a birds-eye-view form the majority of an expert demonstration. Additionally, all of the baselines fail to inpaint the right size of the novel object.
@@ V-C Comparison to Baselines
V-D Data augmentation speed
TABLE III: Breakdown of Minutes Taken for Each Part of Our Method
@@ V-D Data augmentation speed
[…]
We also notice that the majority of the time taken was from the segmentation and background inpainter component (refer to Table III). This module only needs to run once per task as opposed to once per novel object because it deals with erasing the original object in the training videos which is shared across all novel objects. For this reason, we also calculated the time without the training and running of the inpainter, resulting in NeRF-Aug running 4.99 times faster than P2P. In Table II, we show the time taken including the segmentation, training the background NeRF model, and inpainting of the background NeRF (‘w/ inpainter’) and excluding this component (‘w/o inpainter’). We measure all times on a single Nvidia RTX A4000 graphics card.
VI Conclusion

### v3

@@ Abstract
Keywords: Imitation Learning, Novel View Synthesis, Generalization
1 Introduction
@@
Figure 1: When a human provides expert demonstrations for training a behavior cloning model, the model is effective for the object in the demonstration, but will fail for novel objects. We propose NeRF-Aug, where we automatically learn NeRFs for the novel objects and inpaint them in the expert data. With this photorealistic synthetic data, the robot can learn to interact with novel objects.
@@
  • We edit existing expert video demos by removing the training object via in-painting and blend the NeRF render of a novel object into the inpainted image to create synthetic data.
  • We demonstrate effective generalization on five diverse tasks using the generated synthetic dataset for training.
2 Related Works
2.1 Imitation Leaning
@@
2.2 Neural Radiance Fields (NeRFs)
In recent years, there has been a surge in research on NeRFs. [20, 21, 22, 23, 24] are all NeRF approaches that are capable of rendering high resolution images in real time. Other methods focus on the robustness of NeRF models against real-world constraints, such as limited number of images [25], camera pose noise [26], and glare [27]. Finally, other novel-view synthesis methods [28, 29, 30, 31] remove the traditional mlp used to render images with other trainable structures.
Within the context of robotics, NeRFs have been used for robot navigation by creating a 3D map of the environment  [32][33, 34] combine NeRFs and reinforcement learning.  [35, 36] explored using robot arms for creating higher quality NeRF models of objects. Conversely, several works have explored using NeRF models for robot decision making, including [37, 38] which creates accurate depth maps for transparent images using NeRF models, and [39] which creates an object level NeRF and imagines a scene with the object in a different place and queries a vision-language model (VLM) for feedback. Finally, [40] adds noisy viewpoints along a demonstration trajectory and take corrective actions to make a more robust behavior cloning model.
2.3 Synthetic Data in Robotics
@@
3 Preliminaries
Imitation Learning. In imitation learning, we assume a premade list of $N$ expert trajectories $D=\left\{\tau_{i}\right\}_{i=1}^{N}$ where each trajectory consists of state-action pairs $\tau_{i}=\{(s_{k},a_{k})\}_{k=1}^{K}$. In visual imitation learning, we assume access to $n$ cameras where each contributes an image to the state. Thus, our state can be denoted as $s_{k}=(I_{1},I_{2},...,I_{n})$ where $I_{k}$ is an image. In vanilla behavior cloning, the policy $\pi$ is trained offline and learns a mapping between states and actions from our expert dataset by minimizing the loss $\mathcal{L}(\theta)=\mathbb{E}_{(s,a)\sim D}\left[\ell\left(\pi(s),a\right)\right]$ for a given distance metric $\ell$.
@@ 4 Our Approach: NeRF-Aug
4.1 Creating a NeRF model of the Novel Object
Given a novel object for which there exists no training data, we train a NeRF model for this object. Using a gripper mounted camera, we are able to collect a dataset of image-pose pairs $\left\{(I_{k},T_{k})\right\}_{k=1}^{M}$ by moving the gripper to various viewpoints both close and far away from the novel object. We use these images to train a NeRF model for the object. In our experiments, we used NerfStudio’s [48] default Nerfacto model, which refines the camera parameters of all images to denoise measurements.
4.2 Calculating Camera Pose Relative to Object
@@
Next, NeRF-Aug finds the camera position relative to the object. Using the gripper position with respect to the world given by $T_{\text{gripper}}(t)\in\text{SE}(3)$, we are able to find the camera-to-world matrix of the gripper camera. We multiply the gripper position by the offset of the camera from the gripper
@@
If the object is grasped, the relative position of the camera and the object stays the same, denoted by
@@
We also add a small amount of noise at each timestep into the camera-to-object matrix before rendering in the NeRF model. This slightly shifts the object from its original position to simulate a larger degree of grasp positions.
4.3 NeRF Renderings
@@
Figure 3: Inpainting results when replacing the original object with various objects in our training set. Instruct P2P, GenAug, and RoboAgent create images that show a new object, but do not recreate our specific object with appropriate dimensions and texture, causing these methods to fail. Dreambooth has trouble with inpainting complex masks and does not recreate the right shade of colors.
4.4 Combining NeRF Renderings and Original Image
@@
This results in a frame where the original object is removed from the frame and only the background scene remains. From there, we blend the object pixels from the NeRF renderings onto the image where the object was removed. We combine $I_{\text{NeRF}}$ and $I_{\text{no-object}}$ to get a new image with the novel object at the same pose as the original object while still keeping the background of the original image. This is done by computing
@@ 5 Experiments
In our real-world experiments, we use an UR5e robot arm with a Robotiq-85 gripper. The arm is equipped with a Realsense D455 mounted to the front of the gripper (see Fig. 5). We run all experiments on a single Nvidia RTX A4000 GPU. We use the default Nerfacto model from Nerfstudio [48] for training without any hyperparameter tuning aside from stopping the training at 4,000 timesteps for object NeRFs and 20,000 timesteps for background NeRFs.
@@
Table 1: Task success rates for novel objects ($\uparrow$)
@@ 5.1 Baselines
5.2 Real World Tasks
For the Remove Object task, our initial object is a yellow spatula that rests on top of a small frying pan. The robot grasps the spatula and lifts it from the pan then drops it to the left of the pan. For this task we test on 3 novel objects: blue tongs, a red spoon, and a long wooden stick. These objects show how well NeRF-Aug can generalize to different colors and lengths of objects. For this task, we used 30 demonstrations.
For the Place in Chest task, our initial object is a mustard bottle which the robot grasps and places into a wooden chest. We tested on two novel objects: a Swiss Miss box and a paperback book. These objects differ in geometry and texture from the original object. We used 30 demonstrations.
@@ 5.2 Real World Tasks
Table 2: Time To Create New Data in Minutes Including NeRF Model Training ($\downarrow$)
@@
5.3 Comparison to Baselines
As seen in Table 1, our method consistently outperforms other methods on all five tasks. Over the next best method, we achieve an average of 43% increase in the ”Remove Objects” task, a 5% increase in the ”Place In Chest” task, an 80% increase on the ”Cook with Pot” task, and a 50% increase on the ”Lift up High” task. Our method achieved a 85.6% success rate compared to the next best method (Dreambooth) with a 30% success rate.
Which methods synthesize object locations and sizes accurately? GenAug, Instruct Pix-to-Pix, and RoboAgent often output images of an object in either the wrong orientation or position. As seen in fig.  3, we prompted the baselines to inpaint a ”red mug,” only for the output image to consist of a sideways view of a mug that is much larger than the actual mug. In contrast, our method was able to accurately synthesize the mug from a top-down view. Being able to render an object at unconventional viewpoints, such as when an object is very close to the camera or in a birds-eye-view, is a major reason why our method performs better than the baselines (such as being able to lift the mug 90% more often than any other method). Dreambooth seems more effective at synthesizing the correct object viewpoint, but will fail when the original object is significantly different in shape from the old object (see fig. 3 where Dreambooth fails to inpaint a swiss miss box and a book when the original object was a mustard bottle). A wide degree of viewpoints is important because images where the object is very close to the camera or from a birds-eye-view form the majority of an expert demonstration. Additionally, all of the baselines fail to inpaint the right size of the novel object.
@@ 5.3 Comparison to Baselines
5.4 Data augmentation speed
Figure 5: Our setup. We have a Realsense D455 camera, During our tests, we use a large number of background objects such as the two dish racks and the wooden chest to reflect a real world scenario. The right image shows all objects used in experiments. Left side are novel objects, right side are original objects.
Table 3: Breakdown of Minutes Taken for Each Part of Our Method
@@ 5.4 Data augmentation speed
We show that our method is significantly faster than other methods, as seen in Table 2. Our NeRF model is actually able to render at twice the fps as the original video because of our 128 by 128 resolution. Namely, our method is 63.6% faster than the second fastest method (P2P).
We also notice that the majority of the time taken was from the segmentation and background inpainter component (refer to Table 3). This module only needs to run once per task as opposed to once per novel object because it deals with erasing the original object in the training videos which is shared across all novel objects. For this reason, we also calculated the time without the training and running of the inpainter, resulting in NeRF-Aug running 4.99 times faster than P2P. In Table 2, we show the time taken including the segmentation, training the background NeRF model, and inpainting of the background NeRF (‘w/ inpainter’) and excluding this component (‘w/o inpainter’). We measure all times on a single Nvidia RTX A4000 graphics card.
6 Conclusion